# CPRet: A Dataset, Benchmark, and Model for Retrieval in Competitive Programming

### v1

@@ A.6.3 Simplified-to-Full
https://codeforces.com/problemset/problem/1183/F Full problem description Topforces Strikes Back Problem Description One important contest will take place on the most famous programming platform (Topforces) very soon! The authors have a pool of $n$ problems and should choose at most three of them into this contest. The prettiness of the $i$th problem is $a_{i}$ The authors have to compose the most pretty contest (in other words, the cumulative prettinesses of chosen problems should be maximum possible). But there is one important thing in the contest preparation: because of some superstitions of authors, the prettinesses of problems cannot divide each other. In other words, if the prettinesses of chosen problems are $x,y,z$ , then $x$ should be divisible by neither $y$ , nor $z$ , $y$ should be divisible by neither $x$ , nor $z$ and $z$ should be divisible by neither $x$ , nor $y$ If the prettinesses of chosen problems are $x$ and $y$ then neither $x$ should be divisible by $y$ nor $y$ should be divisible by $x$ Any contest composed from one problem is considered good. Your task is to find out the maximum possible total prettiness of the contest composed of at most three problems from the given pool. You have to answer $q$ independent queries. If you are Python programmer, consider using PyPy instead of Python when you submit your code. Input Format The first line of the input contains one integer $q$ ( $1\leq q\leq 2\cdot 10^{5}$ ) — the number of queries. The first line of the query contains one integer $n$ ( $1\leq n\leq 2\cdot 10^{5}$ ) — the number of problems. The second line of the query contains $n$ integers $a_{1},a_{2},\dots,a_{n}$ ( $2\leq a_{i}\leq 2\cdot 10^{5}$ ), where $a_{i}$ is the prettiness of the $i$th problem. It is guaranteed that the sum of $n$ over all queries does not exceed $2\cdot 10^{5}$ Output Format For each query print one integer — the maximum possible cumulative prettiness of the contest composed of at most three problems from the given pool of problems in the query.

### v2

@@ A.8.3 Simplified-to-Full
https://codeforces.com/problemset/problem/1183/F Full problem description Topforces Strikes Back Problem Description One important contest will take place on the most famous programming platform (Topforces) very soon! The authors have a pool of $n$ problems and should choose at most three of them into this contest. The prettiness of the $i$ -th problem is $a_{i}$ . The authors have to compose the most pretty contest (in other words, the cumulative prettinesses of chosen problems should be maximum possible). But there is one important thing in the contest preparation: because of some superstitions of authors, the prettinesses of problems cannot divide each other. In other words, if the prettinesses of chosen problems are $x,y,z$ , then $x$ should be divisible by neither $y$ , nor $z$ , $y$ should be divisible by neither $x$ , nor $z$ and $z$ should be divisible by neither $x$ , nor $y$ . If the prettinesses of chosen problems are $x$ and $y$ then neither $x$ should be divisible by $y$ nor $y$ should be divisible by $x$ . Any contest composed from one problem is considered good. Your task is to find out the maximum possible total prettiness of the contest composed of at most three problems from the given pool. You have to answer $q$ independent queries. If you are Python programmer, consider using PyPy instead of Python when you submit your code. Input Format The first line of the input contains one integer $q$ ( $1\leq q\leq 2\cdot 10^{5}$ ) — the number of queries. The first line of the query contains one integer $n$ ( $1\leq n\leq 2\cdot 10^{5}$ ) — the number of problems. The second line of the query contains $n$ integers $a_{1},a_{2},\dots,a_{n}$ ( $2\leq a_{i}\leq 2\cdot 10^{5}$ ), where $a_{i}$ is the prettiness of the $i$ -th problem. It is guaranteed that the sum of $n$ over all queries does not exceed $2\cdot 10^{5}$ . Output Format For each query print one integer — the maximum possible cumulative prettiness of the contest composed of at most three problems from the given pool of problems in the query.
A.9 Online Demo and Test Cases of CPRetriever
We deployed an open-source competitive programming problem retrieval platform, CPRetriever, on May 21, 2025 (https://cpret.online/). Within the first week, the platform processed nearly 2,000 search queries, and the announcement post on Codeforces received over 250 upvotes, indicating strong community interest.
The platform supports two primary retrieval functionalities: (i) similar problem retrieval, which assists users in expanding problem-solving perspectives and identifying knowledge gaps, and (ii) duplicate problem detection, which aids problem setters in identifying previously seen ideas or solutions.
A.9.1 Duplicate Problem Retrieval in a Recent Contest
We evaluated CPRetriever on the 2025 CCPC National Invitational Contest (Northeast), which featured 12 problems (https://codeforces.com/gym/105924). The system successfully identified six problems with highly similar or identical historical counterparts. Manual inspection of the top three retrievals per query suggests a minimum duplicate rate of 50%, highlighting potential fairness concerns in contest scoring.
Table 9: Detected duplicates in the 2025 CCPC Northeast Contest.
Contest Problem Matched Historical Problem Similarity Level Rank
A. GD Ultimate Rhythm Lab Nowcoder - Xiao Rui Rui’s Sequence Same approach 1
D. Defend the Carrot SPOJ - UOFTBB Almost identical 1
E. Tree Edge Removal Luogu - [JRKSJ R7] Stem Almost identical 1
F. Youthful Oath II Codeforces - 80B Depression Almost identical 1
J. Kingdom: Memories AtCoder - R Walk Almost identical 3
L. Bathhouse Codeforces - 219E Parking Lot Same approach 2
A.9.2 Similar Problem Retrieval: MEX Variants
We further evaluated the system using the classic “interval MEX” problem to retrieve its variants across multiple contests, demonstrating CPRetriever’s utility for idea exploration and knowledge transfer.
Table 10: Top retrieval results for an interval MEX query.
Rank Problem Description
1 Luogu P4137: RMQ Problem / MEX Original problem
2 LOJ 6908: THUPC 2024 Prelim - “Matryoshka” MEX of all subarrays of length $k$, then take MEX
5 AtCoder ABC194E: Mex Min MEX of all subarrays of length $k$, then take minimum
6 Luogu P10032: Mex of Sequence Repeated operations: $a^{\prime}[i]=\mathrm{mex}(a\setminus a[i])$
11 Nowcoder 237670: Classic Problem MEX queries on permutations, optimized to $O(n+m)$
14 Luogu P8087: JROI-5 Interval MEX of all subarrays of length $k$, then take maximum
15 AtCoder ABC290C: Max MEX MEX of all subsequences of length $k$, then take minimum
16 Codeforces 1436E: Complicated Computations MEX of all subarrays, then take MEX again
23 AtCoder ABC330E: Mex and Update Supports element modification or full-array MEX queries
24 Luogu P11837: Making Mexes B Minimum edits to ensure $\mathrm{mex}(a)=i$
These case studies demonstrate that CPRetriever effectively identifies both duplicates and semantically similar problems, supporting practical contest preparation, knowledge expansion, and fairness monitoring in competitive programming environments.
A.10 Figures of data collection and processing pipeline
Problem Sources (Codeforces, AtCoder, Luogu, etc.) Crawling and Data Extraction Match Problems with Accepted Solutions Collect Metadata: Language, Contest Type, Timestamp Multi-language & Multi-format Coverage (ICPC-style, OI-style) Quality Filtering & Data Normalization Annotate with Temporal Information CPRet-PCPCD Dataset (Problems + Solutions + Metadata)
Figure 8: Overall construction pipeline of the CPRet-PCPCD dataset. Problems and accepted solutions are collected from multiple online judges, paired and enriched with metadata, filtered for quality, and annotated with timestamps to support temporally-aware retrieval research.
Discussion Threads & Blog Posts (Codeforces, Luogu) Keyword-based Filtering & LLM Classification 5K Candidate Entries Manual Annotation by Experienced Programmers Three Duplication Levels: Exact / Near / Method Match Clustering Duplicate Problems 700 Duplicate Pairs (Clusters Formed) Split? 30% Clusters $\rightarrow$ Test Set Randomly Select Query & Corpus Problems Add Remaining Codeforces Problems as Distractors Problem-to-Duplicate Retrieval Task SelectSelect
Figure 9: Data collection and processing pipeline for the Problem-to-Duplicate retrieval task. Discussion threads and blogs are filtered and annotated to identify duplicate problems, which are clustered and split into test sets with added distractors for realistic retrieval evaluation.
Original Problems (Codeforces / AtCoder) User-contributed Simplified & Translated Versions (Luogu) Data Crawling Filtering & Quality Control (remove low-quality or MT-based entries) 17K High-quality Simplified–Full Pairs Split? Train Set (7K pairs) Test Set (10K pairs) Simplified-to-Full Retrieval Task (Query = simplified, Corpus = full) 70%30%
Figure 10: Data collection and processing pipeline for the Simplified-to-Full retrieval task. Problems are simplified or translated by Luogu users, then crawled, filtered for quality, and split into training and test sets.